OpenStudy (anonymous):

3/2 times 1/4 ≥ 1/2 (2x+1) ........... Plss answer it ... i cant do this ....

OpenStudy (anonymous):

\[\frac{3}{2} \times \frac{1}{4} \ge \frac{1}{2}\times (2x+1)\]First, what is 3/2 x 1/4?

OpenStudy (anonymous):

is equals to 3/8 and then ?

OpenStudy (anonymous):

Right. Now distribute the 1/2 on the right side to everything inside the parenthesis.

OpenStudy (anonymous):

(2x+1) = 1/2 ????

OpenStudy (anonymous):

i can't understand ?

OpenStudy (anonymous):

Yes you can.\[\frac{1}{2} \times (2x+1)=\frac{1}{2}\times 2x + \frac{1}{2}\times 1\]

OpenStudy (anonymous):

how about 3/8 ?

OpenStudy (anonymous):

We'll come back to that soon. For now, finish simplifying the right side of the equation.

OpenStudy (anonymous):

how ....?

OpenStudy (anonymous):

Well, what is\[\frac{1}{2} \times 2x\]?

OpenStudy (anonymous):

2x X 1 = 3x ... 2x X 2 = 4x

OpenStudy (anonymous):

how to multiply by fraction ?

OpenStudy (anonymous):

If you have\[\frac{a}{b} \times \frac{c}{d}\]when you multiply them together you get\[\frac{a \times c}{b \times d}\]

OpenStudy (anonymous):

And if there is nothing under one of them (say like for the number 2) then you pretend there is a 1 underneath of it.

OpenStudy (anonymous):

the answer it 2x/2

OpenStudy (anonymous):

But what is 2/2?

OpenStudy (anonymous):

how about the x ...?

OpenStudy (anonymous):

\[\frac{2x}{2}=\frac{2}{2}x\]if that helps

OpenStudy (anonymous):

ahhh ... i knew it ..

OpenStudy (anonymous):

and then whats the next process ?

OpenStudy (anonymous):

Well we have\[\frac{3}{8} \ge x+\frac{1}{2}\]now. Subtract 1/2 from both sides.

OpenStudy (anonymous):

2/6

OpenStudy (anonymous):

No.\[\frac{3}{8}-\frac{1}{2}=\frac{3}{8}-\frac{1}{2} \times \frac{4}{4}=\frac{3}{8}-\frac{4}{8}=\frac{3-4}{8}\]You have to make sure that the denominators are the same, and then you add or subtract the numerators.

OpenStudy (anonymous):

-1/8 ???

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ohhh .... thanks ....

OpenStudy (anonymous):

So now you have\[\frac{-1}{8} \ge x\]and that's it!

OpenStudy (anonymous):

owkey ... but the 8 is not negative ?

OpenStudy (anonymous):

Well, if it was, then you would have\[\frac{-1}{-8}=\frac{1 \times -1}{8 \times -1}=\frac{1}{8} \times \frac{-1}{-1}=\frac{1}{8}\]

OpenStudy (anonymous):

owkey ... thanks ... very much ..

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

how about checking ?

OpenStudy (anonymous):

Checking?

OpenStudy (anonymous):

yess ...?

OpenStudy (anonymous):

You mean to make sure you have the right answer? You could take -1/8 and plug it in for x in your original equation. When you simplify both sides they should be the same.

OpenStudy (anonymous):

how ... 3/2 X 1/4 > 1/2 X 2 (-1/8) +1